### v1

# Marginal Metals and Kosterlitz-Thouless Type Phase Transition in Disordered Altermagnets

Abstract
Altermagnetism, a recently discovered magnetic phase characterized by spin-split bands without net magnetization, has emerged as promising platform for novel physics and potential applications. However, its stability against disorder—ubiquitous in real materials—remains poorly understood. Here, we study the electron localization properties of two-dimensional altermagnets subject to disorder. Remarkably, we discover a disorder-driven phase transition from a marginal metallic phase to an insulator, which falls into the Kosterlitz-Thouless class. We demonstrate this by convincing numerical evidence and propose a vortex-antivortex-like spin pair picture for its interpretation. Moreover, we show that the characteristic spin anisotropy of altermagnets persists but gradually fades away across the transition. These changes directly affect the spin-splitting features that are detectable in angle-resolved photoemission spectroscopy and tunneling magnetoresistance. Our findings provide a new perspective on recent experimental observations of altermagnetism in candidate materials.
Introduction.—Altermagnetism is an emerging collinear magnetic phase that uniquely combines advantages of ferromagnetism and antiferromagnetism Naka et al. (2019); Hayami et al. (2019); Šmejkal et al. (2022a, b); Bai et al. (2024); Song et al. (2025). In momentum space, altermagnets (AMs) exhibit highly anisotropic, non-relativistic spin splitting of electronic bands, e.g., manifesting $d$-wave symmetric patterns. From a real-space perspective, this distinctive feature arises because sublattices with opposite spins are related by rotational or mirror symmetry, rather than inversion or translation symmetry Šmejkal et al. (2022a, b). These combined properties renders AMs a fertile platform for exploring novel physics and potential applications in diverse fields such as spintronics, superconductivity, and topological phases González-Hernández, Rafael and Šmejkal, Libor and Výborný, Karel and Yahagi, Yuta and Sinova, Jairo and Jungwirth, Tomáš and Železný, Jakub (2021); Shao et al. (2021); Šmejkal et al. (2022c); Zhu et al. (2023); Zhang et al. (2024); Li and Liu (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Ouassou et al. (2023); Šmejkal et al. (2023); Liu et al. (2024a); Ghorashi et al. (2024); Zhou et al. (2024); Leeb et al. (2024); Sato et al. (2024); Sun et al. (2025); Duan et al. (2025); Gu et al. (2025); Lin et al. (2025); Antonenko et al. (2025); Hu et al. (2025a); Chen et al. (2025). Notably, altermagnetism has been reported in an increasing number of quantum materials Lee et al. (2024); Krempaský et al. (2024); Liu et al. (2024b); Zhu et al. (2025a); Zhang et al. (2025); Ahn et al. (2019); Yuan et al. (2020); Šmejkal et al. (2020); Ma et al. (2021a); Jiang et al. (2025); Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025); Liu et al. (2024c); Das et al. (2024); Xun et al. (2025).
Despite rapidly growing interest in altermagnetism, the impact of disorder, which is inevitably present in real materials, on its stability is barely explored. To date, only a few works have studied the effects of single impurities in AMs Chen et al. (2024); Sukhachov and Linder (2024); Hu et al. (2025b); Maiani and Souto (2025); Gondolf et al. (2025). A crucial open question is thus: What is the fate of the AM phase in presence of disorder? This issue is particularly relevant in light of recent experiments. For example, in confirmed AM candidates at high doping levels, such as $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ Zhang et al. (2025), spin-1/2 dopant atoms may couple opposite electron spins, thereby potentially perturbing the underlying AM order. Moreover, many proposed AM materials, including RuO2 Ahn et al. (2019); Yuan et al. (2020); Šmejkal et al. (2020), $\mathrm{KV_{2}Se_{2}O}$ Ma et al. (2021a); Jiang et al. (2025) and $\mathrm{CrSb}$ Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025), are metallic. In such metallic systems, disorder could significantly alter electronic states and transport properties.
Figure 1: (a) Momentum-resolved local density of states (LDOS) of the $d$-wave AM at disorder strength $W=2t$. (b) Normalized localization length $\lambda/L$ as a function of $W$ for increasing system width $L$. $\lambda$ is calculated on a long ribbon with width $L$ and length of $3\times 10^{6}$. (c) Single-parameter fitting of the correlation length $\xi$ near the KT-type phase transition, extracted from the data in (b). On the insulating side, $\xi$ scales as $\ln\xi\propto|W-W_{c}|^{-1/2}$, where $W_{c}$ denotes the critical disorder strength. Inset: Collapse of the data from (b) into a single curve under finite-size scaling. (d) Phase diagram in the $W$-$t_{J}$ plane. Other parameters are $t_{J}=0.3t$ and $\mu=-2t$.
In this work, we systematically investigate how disorder affects the stability of the AM phase in two dimensions (2D). Strikingly, we find that metallic AMs persist as an exotic marginal metallic state with zero net magnetization over a finite range of disorder strengths, even when scattering between opposite spins becomes significant. Upon further increasing disorder strength, the AMs undergo a distinct metal–insulator phase transition. Through extensive scaling analyses of correlation length and conductance in a representative $d$-wave AM system, we find that this phase transition belongs to the Kosterlitz-Thouless (KT) class. To understand this phase transition, we provide a physical picture based on vortex-antivortex-like spin pairs. Furthermore, we find that the characteristic spin anisotropy of the AM persists but gradually diminishes during the transition. This behavior can be probed experimentally via angle-resolved photoemission spectroscopy (ARPES) and tunneling magnetoresistance. Our results unveil the particular impact of disorder on AMs and suggest disorder as a potential reason for the difficulty in detecting the predicted spin splitting in candidate materials such as RuO2.
Model and general analysis.Altermagnets feature alternating spin splitting in electronic bands, dictated by spin-space group symmetries Šmejkal et al. (2022b); Liu et al. (2022). To elucidate the essential physics, we consider a disordered $d$-wave AM described by the minimal tight-binding Hamiltonian on a square lattice Šmejkal et al. (2022a, b)
$H=$ $-\sum_{\langle{\bf r},{\bf r}^{\prime}\rangle,\sigma,\sigma^{\prime}}\{(t\delta_{\sigma,\sigma^{\prime}}+t_{J}[{\bf e}_{{\bf r}{\bf r}^{\prime}}\cdot\bm{\sigma}]_{\sigma\sigma^{\prime}})c_{{\bf r}\sigma}^{\dagger}c_{{\bf r}^{\prime}\sigma^{\prime}}+h.c.\}$
$-\sum_{{\bf r},\sigma}\mu c_{{\bf r}\sigma}^{\dagger}c_{{\bf r}\sigma}+\sum_{{\bf r},s=0,x,y}c_{{\bf r}\sigma}^{\dagger}w_{{\bf r}}^{s}[\sigma_{s}]_{\sigma\sigma^{\prime}}c_{{\bf r}\sigma^{\prime}},$ (1)
where $c_{{\bf r}\sigma}$ and $c_{{\bf r}\sigma}^{\dagger}$ are electron annihilation and creation operators at position ${\bf r}$ with spin $\sigma\in\{\uparrow,\downarrow\}$. The notation $\langle{\bf r},{\bf r}^{\prime}\rangle$ indicates nearest-neighbor sites. The parameter $t$ represents the normal hopping amplitude and is taken as the energy unit, while $t_{J}$ denotes the spin-dependent alternating hopping strength along different directions associated with the AM order. The unit vector is ${\bf e}_{ij}=\pm{\bf e}_{z}$ for hopping along $x$ ($y$) direction, $\mu$ is the Fermi energy, and $\bm{\sigma}=(\sigma_{x,}\sigma_{y},\sigma_{z})$ are the Pauli matrices for spin.
When disorder is absent, the two spins are decoupled. The system respects $\text{[}C_{2}||C_{4z}]$ spin symmetry, i.e., a fourfold spatial rotation about $z$-axis combined with a spin flip. This symmetry imposes $d$-wave altermagnetism with zero net magnetization. The two spin-polarized bands are given by $E_{\uparrow(\downarrow)}({\bf k})=-2t(\cos k_{x}+\cos k_{y})\mp 2t_{J}(\cos k_{x}-\cos k_{y})-\mu$, where ${\bf k}=(k_{x},k_{y})$ is the 2D momentum and the lattice constant is set to $a=1$. It gives rise to anisotropically spin-split Fermi surfaces, as illustrated in Fig. 1(a).
We consider onsite disorder [the last term in Eq. (1)], which can couple the spin degrees of freedom. Specifically, the $w_{i}^{0}\sigma_{0}$ term represents nonmagnetic disorder, while $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$ correspond to magnetic disorder that flip the spins Šmejkal et al. (2022c); Li2 (a). Here, $w_{i}^{s}$ are random values taken from a uniformly distributed range $[-W/2,W/2]$, with $W$ the disorder strength. Note that this model explicitly breaks time reversal symmetry, placing it in the unitary class (class $A$) in 2D Altland and Zirnbauer (1997); Schnyder et al. (2008). Generally, electronic states in this class will be localized even at infinitesimal disorder strength, except for quantum Hall plateau transitions Wei et al. (1988); Ando (1983). Surprisingly, as we show below, the AM metals exhibit particular extended states even under strong disorder, defying this conventional expectation.
KT-type transition in disordered AM.—To study the stability of AM metals under disorder, we consider a long ribbon geometry with width $L_{y}=L$ and length $L_{x}\gg L$. Whether the states remain delocalized or become localized can be deduced from the localization length under finite-size scaling. We compute the localization length $\lambda$ by means of transfer matrix MacKinnon and Kramer (1983); Kramer and MacKinnon (1993); Yamakage et al. (2013), and define a normalized localization length as $\Lambda\equiv\lambda(L,W)/L$. According to finite-size scaling theory, $\Lambda$ increases with $L$ for metallic states, decreases with $L$ for insulating states, and remains scale-invariant at a critical point MacKinnon and Kramer (1983).
Figure 1(b) shows the scaling behavior of $\Lambda$ in the disordered AM. Remarkably, over a wide range of disorder strengths, $\Lambda$ remains independent of $L$, indicating a marginal metallic state. We refer to this phase as the altermagnetic marginal metal (AMMM). Its emergence suggests that the AM metallic state is resilient to disorder up to a considerable strength, which is crucial for the transport characteristics discussed later. Beyond a critical disorder strength, however, $\Lambda$ decreases monotonically as $L$ grows, indicating that the system is turned into a localized state. It thus realizes a disorder-driven metal-insulator transition. The appearance of the scale-invariant AMMM phase implies that this transition belongs to KT class.
To verify the nature of the transition, we analyze the scaling behavior of the correlation length $\xi$ (corresponding to $\lambda$ in the $L\rightarrow\infty$ limit) on the insulating side. In KT-type metal-insulator transitions, $\xi$ diverges exponentially near the critical disorder strength $W_{c}$, following a universal form $\xi\propto\exp[b/\sqrt{W-W_{c}}]$, where $b$ is a constant parameter Xie et al. (1998); Wang et al. (2015); Chen et al. (2019). Indeed, as shown in Fig. 1(c), our numerical data nicely match the scaling form
$\ln\xi=b|W-W_{c}|^{-1/2}$ (2)
with fitting parameters $b=9.24\pm 0.21$ and $W_{c}=4.13t\pm 0.08t$. Moreover, all finite-size scaling curves for different $L$ collapse to a single curve [inset of Fig. 1(c)]. These results strongly support that the metal-insulator transition is of KT type  [^1] For KT transition in 2D XY model, the correlation length scales as $\xi(T)\propto\exp[b/\sqrt{T-T_{c}}]$ above $T_{C}$ with $T$ the temperature.. Additionally, by repeating the above analysis for varying values of $t_{J}$, we map out the phase diagram in the $t_{J}$-$W$ plane, shown in Fig. 1(d). The diagram shows that $W_{c}$ increases with $t_{J}$, indicating enhanced robustness of the metallic phase by stronger altermagnetism.
Figure 2: (a) Disorder-averaged logarithmic conductance $\langle\ln g\rangle$ as a function of $W$, with each point averaged over 3000 disorder configurations. (b) Scaling function $\beta\equiv\frac{d\langle\ln g\rangle}{d\ln L}$ extracted from (a). Other parameters are as the same as Fig. 1.
Scaling features of conductance.—Further evidence for the KT-type phase transition can be obtained from transport properties. We consider a square AM sample of length $L$, connected to two conducting leads. The two-terminal conductance $g_{L}$ can be calculated using the Landauer-Büttiker formalism as $g_{L}=\mathrm{Tr}[TT^{\dagger}]$ Datta (1995), where $T$ is the transmission matrix obtained via the recursive Green’s function method MacKinnon (1985); Sancho et al. (1985). To extract the intrinsic transport nature, we subtract the contact resistance between the AM sample and the leads. Thus, the bulk conductance $g$ is defined as $1/g=1/g_{L}-1/N_{C}$, where $N_{C}$ is the number of conducting channels at the Fermi energy Slevin et al. (2001); Zhang et al. (2009).
Figure 2(a) shows the scaling of the disorder-averaged conductance $\langle\ln g\rangle$. Remarkably, in the AMMM phase ($W<W_{c}$), curves for different $L$ merge together, indicating a finite conductance even in the thermodynamic limit. In contrast, on the insulating side ($W>W_{c}$), $\langle\ln g\rangle$ decreases monotonically with increasing $L$, implying that all states become localized. These trends are consistent with the localization length analysis. To further characterize the transition, we compute the scaling function $\beta\equiv\frac{d\langle\ln g\rangle}{d\ln L}$ Abrahams et al. (1979). Typically, a positive (negative) $\beta$ signals delocalized (localized) states. As shown Fig. 2(b), we find $\beta=0$ in the AMMM region, indicating delocalized states irrespective of system size. For strong $W(>W_{c})$, $\beta<0$ as expected for the insulating phase. The analysis of $\beta$ supports the KT metal-insulator transition in disordered AMs.
Spectral statistics.—For a comprehensive understanding of disordered AMs, we examine the spectral statistics using level spacing ratio (LSR). The LSR is defined as $r_{n}=\frac{\mathrm{min}(s_{n},s_{n-1})}{\mathrm{max}(s_{n},s_{n-1})}$, where $s_{n}\equiv E_{n+1}-E_{n}$ is the spacing between adjacent energy levels $E_{n}$ and $E_{n+1}$ Oganesyan and Huse (2007). Figure S3(a) plots disorder-averaged LSR $\langle r\rangle$ against disorder strength $W$. For a finite range of $W<W_{c}$, $\langle r\rangle$ remains nearly constant at $\langle r\rangle\approx 0.6$, consistent with the metallic nature of AMMMs ($\langle r\rangle_{M}=0.6$). As $W$ exceeds $W_{c}$, $\langle r\rangle$ decreases smoothly toward $\langle r\rangle_{I}=0.386$, the universal value characteristic of localized insulating states Atas et al. (2013). The probability distributions $P(r)$ at representative $W$ are shown in Fig. S3(b). For weak disorder ($W=3t$), $P(r)$ closely follows the Gaussian unitary ensemble (GUE) form $P_{\text{GUE}}(r)=\frac{81\sqrt{3}}{2\pi}\frac{(r+r^{2})^{2}}{(1+r+r^{2})^{4}}$ Atas et al. (2013), which is typical of extended states. For strong disorder ($W=14t$), $P(r)$ instead matches the Poisson distribution $P_{p}(r)=\frac{2}{(1+r)^{2}}$, indicative of uncorrelated, localized states. In the SM, we also analyze inverse participation ratio statistics Li2 (b). These results further confirm the existence of a robust AMMM phase and the disorder-induced metal–insulator transition.
Figure 3: (a) Averaged LSR $\langle r\rangle$ as a function of $W$. (b) Distribution of LSR in the metallic and insulating limits.Other parameters are as the same as Fig. 1.
Phenomenological picture.—We present a phenomenological argument to interpret the KT-type phase transition in disordered AMs. In the 2D XY model, the KT transition arises from the unbinding of vortex-antivortex pairs as temperature increases Kosterlitz and Thouless (1973). Analogously, the spin-up and spin-down degrees of freedom in disordered AMs can be viewed as vortex-like and antivortex-like excitations, respectively. Weak magnetic disorder, described by the terms $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$, couples spin-up and spin-down states (quantized along $\sigma_{z}$), effectively forming spin-paired states ($\bm{\uparrow}\leftrightsquigarrow\bm{\downarrow}$) akin to bound vortex-antivortex pairs thermally activated in the XY model. Crucially, the $[C_{2}||C_{4z}]$ spin symmetry (equivalently, the $C_{4}\mathcal{T}$ symmetry that combines four-fold rotation and time-reversal symmetries) protects the marginal metallic phase Wang et al. (2024). Indeed, as shown in Fig. 1(d), the critical disorder strength $W_{c}$ increases with the AM strength $t_{J}$. In absence of AM ($t_{J}=0$) or in conventional antiferromagnets, no marginal metal phase emerges, underscoring the essential role of altermagnetism. When the disorder strength exceeds $W_{c}$, magnetic scattering between opposite spins becomes so strong that the spin pairs are fully depaired, resembling the breakdown of vortex-antivortex pairs in the XY model. The system then enters an insulating state. In this way, the AM undergoes a KT-type phase transition driven by disorder.
Figure 4: (a) Spin-resolved LDOS $A_{\sigma}({\bf k})$ (left and middle) and total LDOS $A({\bf k})=A_{\uparrow}({\bf k})+A_{\downarrow}({\bf k})$ (right) in momentum space for disorder strength $W=t$. (b) The same as (a) but for $W=6t$. Other parameters are $t_{J}=0.3t,$ $\mu=-2t$, and $\eta=10^{-3}t$.
Mitigation of spin anisotropy by disorder.—Disorder significantly affects the characteristic spin anisotropy of AMs, namely, the alternating spin splitting in momentum space. To illustrate this, we track the evolution of spin anisotropy during the phase transition by analyzing spin-resolved LDOS. The LDOS is calculated from the disorder-averaged effective Hamiltonian $\langle H({\bf k})\rangle$ via
$A_{\sigma}(E,{\bf k})=-\frac{1}{\pi}\mathrm{Im}[E+i\eta-\langle H({\bf k})\rangle]_{\sigma\sigma}^{-1},$ (3)
where $\eta$ is a small spectral broadening Li2 (b). As shown in Fig. 4(a), the spin anisotropy remains prominent in the AMMM phase at weak disorder: spin-up and spin-down Fermi surfaces exhibit distinct orientations, confirming the persistence of spin-splitting. As disorder strength $W$ increases, the bands broaden due to disorder scattering, and the Fermi-surface splitting gets increasingly blurred [see Figs. 4(a,b) for comparison]. In this regime, the anisotropic $d$-wave AM gradually evolves towards an isotropic pattern. This transition process is similar to the disorder-driven suppression of $d$-wave superconducting pairing Balatsky et al. (2006); Spivak et al. (2009). When $W$ exceeds $W_{c}$, spin anisotropy rapidly fades away [Fig. 4(b)]. In this strong-disorder regime, spin-up and spin-down states strongly scatter into each other, and the system eventually enters an insulating phase. Note that throughout this evolution, the system maintains zero net magnetization, and spin remains a good quantum number after disorder average.
The disorder-induced degradation of spin anisotropy in AMs may shed light on recent experiments on candidate materials such as $\mathrm{RuO_{2}}$ and $\mathrm{MnF_{2}}$ Fedchenko et al. (2024); Lin et al. (2024); Liu et al. (2024d); Morano et al. (2025). In $\mathrm{RuO_{2}}$, while some ARPES measurements report spin-splitting features Fedchenko et al. (2024); Lin et al. (2024), others yield conflicting results Liu et al. (2024d). Previous studies attempt to attribute this difficulty to potential Ru vacancies Smolyanyuk et al. (2024) or insufficient electron interactions Qian et al. (2025). Our findings suggest an alternative explanation based on variations in disorder levels of different samples. In samples with weak disorder, spin anisotropy is preserved, enabling detection of spin-split bands. In contrast, in samples with sufficiently strong disorder, spin anisotropy is fully suppressed, rendering the spin-splitting features unobservable.
Figure 5: (a) Schematic of the tunneling junction composed of two AM layers of length $L_{x}$ separated by an insulating barrier. The Néel vectors of two AM layers are configured to be either parallel (P) or antiparallel (AP). The system size is $L_{x}=10a$ and $L_{y}=40a$. (b) Tunneling conductances $G_{P}$ and $G_{AP}$ as functions of energy $E$ for the parallel and antiparallel configurations at $W=0$. (c) $G_{P}$ and $G_{AP}$ as functions of $W$ at different $E$. (d) TMR as a function of $W$. Inset: Energy spectrum of an AM ribbon of width $L_{y}=40$ along $x$ direction. Other parameters are as the same as Fig. 1.
Tunneling magnetoresistance in disordered AMs.The spin splitting affected by disorder in AMs can also be detected via tunneling magnetoresistance in tunnel junctions. We consider a junction composed of two AM layers with either parallel or antiparallel Néel vectors, separated by a tunnel barrier, as illustrated in Fig. 5(a). Recent studies have shown that tunneling magnetoresistance naturally arises in such junctions owing to the intrinsic spin splitting Shao et al. (2021); Šmejkal et al. (2022c); Liu et al. (2024e); Noh et al. (2025). Within the energy range of the band width, we calculate the tunneling conductance using $G=\frac{e^{2}}{h}\mathrm{Tr}[TT^{\dagger}]$, where $T$ is the transmission matrix across the junction. Due to the spin anisotropy, the conductance for the parallel configuration $G_{P}$ is generally larger than that of the antiparallel configuration $G_{AP}$, i.e., $G_{P}>G_{AP}$. This feature is confirmed in Fig. 5(b), which shows $G_{P}$ and $G_{AP}$ as functions of energy. The peaks in $G_{P}(E)$ originate from tunneling resonances in the junction.
We introduce disorder into the two AM layers and calculate the tunneling conductances as functions of disorder strength $W$ [Fig. 5(c)]. We find that a substantial conductance difference between $G_{P}$ and $G_{AP}$ persists over a wide range of $W$, extending even to values $W>W_{c}$. This reflects the fact that spin anisotropy remains robust against weak disorder in the AMMM phase. However, for strong disorder ($W\gtrsim 10t$), $G_{P}$ and $G_{AP}$ fully converge. This signals the complete suppression of spin anisotropy due to disorder scattering. Interestingly, both $G_{P}$ and $G_{AP}$ exhibit nonmonotonic behavior with increasing $W$, showing an enhancement over a finite range of $W$ before eventually merging and decreasing. This increase may be attributed to the fact that AM order reduces while the states are not fully localized. To quantify the sensitivity of tunneling to the AM bilayer configuration, we further compute the tunneling magnetoresistance $\text{TMR}=\frac{G_{P}-G_{AP}}{G_{AP}}$. As shown in Fig. 5(d), TMR depends on energy $E$. At $E=2t$, it decreases monotonically to zero, while at $E=0$, it first increases to a maximum before it eventually drops to zero. The latter case suggests that within a specific range of disorder strengths, the tunneling magnetoresistance can be anomalously enhanced.
Conclusion and discussion.—In summary, we have shown that 2D metallic AMs persist as a marginal metal over a finite range of disorder strengths and undergo a KT metal-insulator transition as disorder increases further. We have also provided a phenomenological interpretation of this transition based on vortex-antivortex-like spin pairs. During the transition, the spin anisotropy of the AM gradually weakens, and the tunneling magnetoresistance becomes indistinguishable. These predictions are experimentally accessible and may help explain conflicting experiments on AM candidate materials.
The key ingredient underlying our results is the interplay between AM order and magnetic disorder. Notably, neither this marginal metallic phase nor the KT-type metal-insulator transition appears in conventional antiferromagnets and nonmagnetic electron gases, where the electronic bands are spin-degenerate. Recently, many 2D AMs have been proposed and observed experimentally, including Mn2PSe with broken inversion symmetry Liu et al. (2025), $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ with layered structures Zhang et al. (2025); Jiang et al. (2025), ferroelectric systems Zhu et al. (2025b), monolayer Ma et al. (2021b) and twisted magnetic bilayers Liu et al. (2024b). Although our study focuses on representative $d$-wave AMs, we expect our conclusions to extend to other symmetry types, such as $g$- and $i$-wave AMs.
We thank Lunhui Hu, K. Mæland, E. Petermann, and Jinsong Xu for helpful discussions. This work was supported by the DFG (SFB 1170 ToCoTronics), and the Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147 (Project-Id 390858490). H.G. acknowledge support from the NSFC grant No. 12074022 and the BNL-CMP open research fund under Grant No. 2024BNL-CMPKF023. S.B.Z was supported by the start-up fund at HFNL, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302801).
Supplemental materials of “Marginal Metals and Kosterlitz-Thouless Type Phase Transition in Disordered Altermagnets”
Appendix S1 Localization length and transfer matrix method
In this section, we outline the transfer matrix method based on the lattice model to obtain the localization length of the system (MacKinnon and Kramer, 1983; Kramer and MacKinnon, 1993). Specifically, we consider a long ribbon geometry along $x$-direction with width $L_{y}=L$ and length $L_{x}\gg L$. In practice, we take $L_{x}=3\times 10^{6}$. Periodic boundary conditions are taken along $y$-direction. We partition the ribbon lattice to $L_{x}$ layers along $x$-direction. Focusing on three consecutive layers at $x=n-1$, $n$, and $n+1$, respectively, their wave function amplitudes can be connected by
$\left(\begin{array}[]{c}\psi_{n+1}\\ \psi_{n}\end{array}\right)=T_{n}\left(\begin{array}[]{c}\psi_{n+1}\\ \psi_{n}\end{array}\right),$ (S1.1)
where the transfer matrix is
$T_{n}=\left(\begin{array}[]{cc}-H_{n,n+1}^{-1}(E\bm{I}-H_{n})&-H_{n,n+1}^{-1}H_{n,n-1}\\ \bm{I}&0\end{array}\right).$ (S1.2)
Here, $H_{n}$ is the $2L\times 2L$ Hamiltonian of the $n$-th layer, $H_{n,n+1}$ is the $2L\times 2L$ interlayer hopping matrix between the $n$-th and $(n+1)$-th layers, and $\bm{I}$ is a $2L\times 2L$ identity matrix.
To extract the localization length, we calculate the transfer matrix product across the whole ribbon,
$O_{M}=\prod_{n=1}^{L_{x}}T_{n},$ (S1.3)
and construct
$P=O_{M}O_{M}^{\dagger}.$
According to the Oseledec theorem, positive eigenvalues $\nu_{i}$ of the matrix $P$ exist, and its logarithm defines the Lyapunov exponents in the limit $L_{x}\rightarrow\infty$ as
$\gamma_{i}=\lim_{L_{x}\rightarrow\infty}\frac{\ln\nu_{i}}{2L_{x}},$ (S1.4)
The quasi-1D localization length is given by the inverse of the smallest Lyapunov exponent
$\lambda=\frac{1}{\min(\gamma_{i})}.$ (S1.5)
In Fig. 1 of the main text, we show the results for $t_{J}=0.3t$. Here, in Fig. S1(a), we further present the localization length for $t_{J}=2.0$ as a representative case of strong altermagnetic strength $t_{J}$. Clearly, the critical disorder strength $W_{c}$ shifts to higher values, reflecting the enhanced resilience of altermagnetic order. By fitting the scaling behavior to the form $\xi\propto\exp[b/\sqrt{W-W_{c}}]$ near $W_{c}$ [Fig. S1(b)], we find $W_{c}=11.27t\pm 0.33t$. We find that all data converge to a single curve under this scaling formula.
Figure S1: Kosterlitz-Thouless (KT) metal-insulator transition for $t_{J}=2.0t$. (a) Normalized localization length $\lambda/L$ as a function of $W$ for increasing $L$. (b) Single-parameter fitting of the correlation length $\xi$ near the KT-type transition, extracted from the data in (a). It gives the critical disorder strength $W_{c}=11.27t\pm 0.33t$. (c) Collapse of the data from (a) into a single curve under finite-size scaling. Other parameters are $t_{J}=2t$ and $\mu=-2t$.
Appendix S2 Disorder averaged effective Hamiltonian
In this section, we describe in details how to obtain the disorder-averaged effective Hamiltonian $\langle H\rangle$. Consider the altermagnetic metal on a $L\times L$ square lattice with periodic boundary conditions in both $x$ and $y$ directions. For each disorder configuration, we find the retarded Green’s function $G^{R}({\bf r},{\bf r}^{\prime};\omega)$ in the tight-binding basis as
$G^{R}({\bf r},{\bf r}^{\prime};\omega)$ $=\langle{\bf r}|(\omega+i\eta-H_{\text{rand}})^{-1}|{\bf r}^{\prime}\rangle,$ (S2.1)
where $\omega$ is the frequency, $\eta$ is an infinitesimal positive number, and $H_{\text{rand}}$ is the lattice Hamiltonian including disorder. Averaging over many disorder configurations, translation invariance is effectively restored. Thus, the disorder-averaged Green’s function takes the form
$G_{\mathrm{avg}}^{R}({\bf r}-{\bf r}^{\prime},\omega)$ $=\langle G^{R}({\bf r},{\bf r}^{\prime};\omega)\rangle_{\mathrm{dis}}.$ (S2.2)
It depends only on the position difference ${\bf r}-{\bf r}^{\prime}$. Next, we next Fourier transform the disorder-averaged Green’s function into momentum space,
$G^{R}({\bf k},\omega)$ $=\int d({\bf r}-{\bf r}^{\prime})G_{\mathrm{avg}}^{R}({\bf r}-{\bf r}^{\prime},\omega)e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}.$ (S2.3)
Finally, the disorder-average effective Hamiltonian is then given by
$\langle H({\bf k})\rangle=-[G^{R}({\bf k},\omega=0)]^{-1}.$ (S2.4)
The spin-resolved density of states then follows from
$A_{\sigma}(E,{\bf k})=-\frac{1}{\pi}\mathrm{Im}[E+i\eta-\langle H({\bf k})\rangle]_{\sigma\sigma}^{-1}.$ (S2.5)
Appendix S3 Scaling behavior under different disorder types
In Fig. S2(a), we consider on-site nonmagnetic (spin-independent) disorder, i.e. $w_{i}^{0}\sigma_{0}$. Since this type of disorder does not mix opposite spins, the altermagnets (AMs) decouple into two independent spin sectors. From Fig. S2(a), we find that the normalized localization length always decreases with system size $L$. This indicates an immediate transition into insulating phases and the absence of Kosterlitz-Thouless (KT) type phase transition, consistent with conventional 2D electron behavior.
This behavior changes completely if we consider magnetic disorder, i.e., $(w_{i}^{x}\sigma_{x}+w_{i}^{y}\sigma_{y}+w_{i}^{z}\sigma_{z})$. Here, the diagonal type $w_{i}^{z}\sigma_{z}$ is included for comparison, although it does not flip spins. As shown in Fig. S2(b), $\Lambda=\lambda/L$ remains independent of $L$ within a wide range of disorder strengths, indicating the existence of a marginal metallic state. Beyond the critical value of disorder strength, $\Lambda$ decreases with $L$, indicating the KT-type metal-insulator phase transition. Comparing these results, we conclude that the magnetic scattering between opposite spins is essential for the occurrence of KT-type phase transitions in 2D AM metals.
Figure S2: (a) Normalized localization length $\lambda/L$ as a function of $W$ for nonmagnetic disorder. (b) Normalized localization length $\lambda/L$ as a function of $W$ for magnetic disorder. Other parameters are $t_{J}=0.5t$ and $\mu=-2t$.
Appendix S4 Spectrum statistics in disordered altermagnetism from inverse participation ratio
In this section, we analyze the statistics from the eigenstates of the disordered AMs. To this end, we employ the inverse participation ratio (IPR) defined as (Li et al., 2017; Pixley et al., 2015)
$I_{n}=\sum_{i,\sigma}|\psi_{n}(i,\sigma)|^{4},$ (S4.1)
where the summation runs over all sites and spins. The disorder-averaged IPR $\langle I\rangle$ takes nearly zero values at weak disorder strength, consistent with the altermagnetic marginal metals discussed in the main text [Fig. S3(a)]. It becomes finite when the system enters an insulating phase.
The IPR scaling defines a fractal dimension $d_{2}$ via $\langle I\rangle\propto L^{-d_{2}}$. From Fig. S3(b), we find $d_{2}=1.84$ in the altermagnetic marginal metals, which is close to the value $d_{2}^{M}=2$ for ideal metals. For the insulating state at $W=14t$, we find $d_{2}=0.14$, consistent with the value $d_{2}^{I}=0$ for insulators. Between these two limits, a smooth transition is observed. Spectral rigidity, quantified by the spectrum compressibility
$\chi=\frac{\langle N^{2}\rangle-\langle N\rangle^{2}}{\langle N\rangle}$ (S4.2)
also reflects the localization and delocalization behavior. Here, $\langle N\rangle$ is the disorder-averaged number of energy levels within an energy window. In 2D, $\chi$ is conjectured to be related to $d_{2}$ by $\chi=(2-d_{2})/4$. As shown in Fig. S3(c), we find $\chi=0.035$ at $W=3t$ and $\chi=0.44$ at $W=14t$, both in excellent agreement with the values inferred from $d_{2}$. Thus, the conjectured relation is confirmed in both the altermagnetic marginal metals and insulating states.
Figure S3: (a) Disorder-averaged IPR as a function of disorder strength $W$. (b) Scaling behavior of IPR for different $W$. (c) Spectrum fluctuation $\Sigma_{2}\equiv\langle N^{2}\rangle-\langle N\rangle^{2}$ vs $\langle N\rangle$ for different energy windows. Other parameters for all plots are $t_{J}=0.3t$, and $\mu=-2t$.

### v2

@@
 However, its stability against disorder—ubiquitous in real materials—remains poorly understood.
-Here, we study the electron localization properties of two-dimensional altermagnets subject to disorder.
+Here, we study the electron localization properties of two-dimensional $d$-wave altermagnets subject to disorder.
 Remarkably, we discover a disorder-driven phase transition from a marginal metallic phase to an insulator, which falls into the Kosterlitz-Thouless class.
-We demonstrate this by convincing numerical evidence and propose a vortex-antivortex-like spin pair picture for its interpretation.
+We demonstrate this by strong numerical evidence and propose an interpretation in terms of vortex-antivortex pairs in the disorder-induced local in-plane spin magnetization.
 Moreover, we show that the characteristic spin anisotropy of altermagnets persists but gradually fades away across the transition.
-These changes directly affect the spin-splitting features that are detectable in angle-resolved photoemission spectroscopy and tunneling magnetoresistance.
+These changes directly affect the spin-splitting features that are detectable in angle-resolved photoemission spectroscopy and tunneling magnetoconductance.
 Our findings provide a new perspective on recent experimental observations of altermagnetism in candidate materials.
-Introduction.—Altermagnetism is an emerging collinear magnetic phase that uniquely combines advantages of ferromagnetism and antiferromagnetism Naka et al.
+Introduction.—Altermagnetism is an emerging collinear magnetic phase that uniquely combines advantages of ferromagnetism and antiferromagnetism Šmejkal et al.
+(2020); Naka et al.
+(2019); Ahn et al.
 (2019); Hayami et al.
+(2019); Yuan et al.
 (2020); Šmejkal et al.
@@
 (2022a, b).
-These combined properties renders AMs a fertile platform for exploring novel physics and potential applications in diverse fields such as spintronics, superconductivity, and topological phases González-Hernández, Rafael and Šmejkal, Libor and Výborný, Karel and Yahagi, Yuta and Sinova, Jairo and Jungwirth, Tomáš and Železný, Jakub (2021); Shao et al.
+These combined properties renders AMs a fertile platform for exploring novel physics and potential applications in diverse fields such as spintronics, superconductivity, and topological phases R.
+González-Hernández, L.
+Šmejkal, Karel Výborný, Y.
+Yahagi, J.
+Sinova, T.
+Jungwirth, and J.
+Železný (2021); Shao et al.
 (2021); Šmejkal et al.
@@
 (2025a); Zhang et al.
-(2025); Ahn et al.
+(2025a); Ahn et al.
 (2019); Yuan et al.
@@
 (2021a); Jiang et al.
-(2025); Reimers et al.
+(2025); Zhang et al.
+(2025b); Reimers et al.
 (2024); Ding et al.
@@
 For example, in confirmed AM candidates at high doping levels, such as $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ Zhang et al.
-(2025), spin-1/2 dopant atoms may couple opposite electron spins, thereby potentially perturbing the underlying AM order.
-Moreover, many proposed AM materials, including RuO2 Ahn et al.
-(2019); Yuan et al.
-(2020); Šmejkal et al.
+(2025a), spin-1/2 dopant atoms may couple opposite electron spins, thereby potentially perturbing the underlying AM order.
+Moreover, many proposed AM materials, including RuO2 Šmejkal et al.
+(2020); Ahn et al.
 (2019), $\mathrm{KV_{2}Se_{2}O}$ Ma et al.
 (2021a); Jiang et al.
-(2025) and $\mathrm{CrSb}$ Reimers et al.
+(2025); Zhang et al.
+(2025a) and $\mathrm{CrSb}$ Reimers et al.
 (2024); Ding et al.
@@
 (b) Normalized localization length $\lambda/L$ as a function of $W$ for increasing system width $L$.
-$\lambda$ is calculated on a long ribbon with width $L$ and length of $3\times 10^{6}$.
+$\lambda$ is calculated on a long ribbon with width $L$ and length of $3\times 10^{6}a$.
 (c) Single-parameter fitting of the correlation length $\xi$ near the KT-type phase transition, extracted from the data in (b).
@@
 In this work, we systematically investigate how disorder affects the stability of the AM phase in two dimensions (2D).
-Strikingly, we find that metallic AMs persist as an exotic marginal metallic state with zero net magnetization over a finite range of disorder strengths, even when scattering between opposite spins becomes significant.
-Upon further increasing disorder strength, the AMs undergo a distinct metal–insulator phase transition.
+Strikingly, we find that metallic AMs persist as an exotic marginal metal with zero net magnetization over a finite range of disorder strengths, even when opposite-spin scattering becomes significant.
+Beyond a critical disorder strength, the system undergoes a distinct metal–insulator transition.
 Through extensive scaling analyses of correlation length and conductance in a representative $d$-wave AM system, we find that this phase transition belongs to the Kosterlitz-Thouless (KT) class.
-To understand this phase transition, we provide a physical picture based on vortex-antivortex-like spin pairs.
-Furthermore, we find that the characteristic spin anisotropy of the AM persists but gradually diminishes during the transition.
-This behavior can be probed experimentally via angle-resolved photoemission spectroscopy (ARPES) and tunneling magnetoresistance.
-Our results unveil the particular impact of disorder on AMs and suggest disorder as a potential reason for the difficulty in detecting the predicted spin splitting in candidate materials such as RuO2.
-Model and general analysis.Altermagnets feature alternating spin splitting in electronic bands, dictated by spin-space group symmetries Šmejkal et al.
+To understand this transition, we propose a physical picture based on vortex-antivortex pairs in the induced local in-plane spin magnetization.
+Furthermore, the characteristic spin anisotropy persists at weak disorder but gradually diminishes across the transition.
+This behavior can be probed via angle-resolved photoemission spectroscopy (ARPES) and tunneling magnetoconductance.
+The results hold for both uncorrelated and correlated disorder and generalize to $g$-wave AMs and other lattices, for instance, the Lieb lattice.
+Our findings underscore the particular impact of disorder on AMs and suggest disorder as a potential reason for the difficulty in detecting the predicted spin splitting in candidate materials such as RuO2.
+Model and general analysis.Altermagnets feature alternating spin splitting in electronic bands, related to spin-space group symmetries Šmejkal et al.
 (2022b); Liu et al.
 (2022).
-To elucidate the essential physics, we consider a disordered $d$-wave AM described by the minimal tight-binding Hamiltonian on a square lattice Šmejkal et al.
+For concreteness, we consider a disordered $d$-wave AM described by the minimal tight-binding Hamiltonian on a square lattice Šmejkal et al.
 (2022a, b)
@@
 When disorder is absent, the two spins are decoupled.
-The system respects $\text{[}C_{2}||C_{4z}]$ spin symmetry, i.e., a fourfold spatial rotation about $z$-axis combined with a spin flip.
-This symmetry imposes $d$-wave altermagnetism with zero net magnetization.
+The system respects $\text{[}C_{2}||C_{4z}]$ spin symmetry, i.e., a fourfold spatial rotation about the $z$-axis combined with a spin flip.
+This symmetry imposes $d$-wave altermagnets described by $H_{0}({\bf k})=-(2t\cos k_{x}+2t\cos k_{y}+\mu)\sigma_{0}-2t_{J}(\cos k_{x}-\cos k_{y})\sigma_{z}$ with zero net magnetization.
 The two spin-polarized bands are given by $E_{\uparrow(\downarrow)}({\bf k})=-2t(\cos k_{x}+\cos k_{y})\mp 2t_{J}(\cos k_{x}-\cos k_{y})-\mu$, where ${\bf k}=(k_{x},k_{y})$ is the 2D momentum and the lattice constant is set to $a=1$.
@@
 (1)], which can couple the spin degrees of freedom.
-Specifically, the $w_{i}^{0}\sigma_{0}$ term represents nonmagnetic disorder, while $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$ correspond to magnetic disorder that flip the spins Šmejkal et al.
-(2022c); Li2 (a).
-Here, $w_{i}^{s}$ are random values taken from a uniformly distributed range $[-W/2,W/2]$, with $W$ the disorder strength.
+Specifically, the $w_{{\bf r}}^{0}\sigma_{0}$ term represents nonmagnetic disorder, while $w_{{\bf r}}^{x}\sigma_{x}$ and $w_{{\bf r}}^{y}\sigma_{y}$ correspond to magnetic disorder that flip the spins Šmejkal et al.
+(2022c) [^1] Note that the main results we discuss will remain if the disorder of form $w_{{\bf r}}^{z}\sigma_{z}$ is added.
+This type of disorder does not flip the spin.
+Here, $w_{{\bf r}}^{s}$ are random values taken from a uniformly distributed range [W/2,W/2][-W/2,W/2], with $W$ the disorder strength.
 Note that this model explicitly breaks time reversal symmetry, placing it in the unitary class (class $A$) in 2D Altland and Zirnbauer (1997); Schnyder et al.
@@
 Surprisingly, as we show below, the AM metals exhibit particular extended states even under strong disorder, defying this conventional expectation.
-KT-type transition in disordered AM.—To study the stability of AM metals under disorder, we consider a long ribbon geometry with width $L_{y}=L$ and length $L_{x}\gg L$.
+KT-type transition in disordered AM.—To study the stability of AM metals in presence of disorder, we consider a long ribbon geometry with width $L_{y}=L$ and length $L_{x}\gg L$.
 Whether the states remain delocalized or become localized can be deduced from the localization length under finite-size scaling.
-We compute the localization length $\lambda$ by means of transfer matrix MacKinnon and Kramer (1983); Kramer and MacKinnon (1993); Yamakage et al.
+We compute the localization length $\lambda$ by means of the transfer matrix MacKinnon and Kramer (1983); Kramer and MacKinnon (1993); Yamakage et al.
 (2013), and define a normalized localization length as $\Lambda\equiv\lambda(L,W)/L$.
@@
 Its emergence suggests that the AM metallic state is resilient to disorder up to a considerable strength, which is crucial for the transport characteristics discussed later.
-Beyond a critical disorder strength, however, $\Lambda$ decreases monotonically as $L$ grows, indicating that the system is turned into a localized state.
+Beyond a critical disorder strength, however, $\Lambda$ decreases monotonically as $L$ grows, indicating that the system is turned into an insulator.
 It thus realizes a disorder-driven metal-insulator transition.
-The appearance of the scale-invariant AMMM phase implies that this transition belongs to KT class.
+The appearance of the scale-invariant AMMM phase implies that this transition belongs to the KT class.
 To verify the nature of the transition, we analyze the scaling behavior of the correlation length $\xi$ (corresponding to $\lambda$ in the $L\rightarrow\infty$ limit) on the insulating side.
@@
 These results strongly support that the metal-insulator transition is of KT type  [^2] For KT transition in 2D XY model, the correlation length scales as $\xi(T)\propto\exp[b/\sqrt{T-T_{c}}]$ above $T_{C}$ with $T$ the temperature..
-Additionally, by repeating the above analysis for varying values of $t_{J}$, we map out the phase diagram in the $t_{J}$-$W$ plane, shown in Fig.
+Additionally, by performing the above analysis for different values of $t_{J}$, we map out the phase diagram in the $t_{J}$-$W$ plane, shown in Fig.
 1(d).
@@
 The analysis of $\beta$ supports the KT metal-insulator transition in disordered AMs.
-Spectral statistics.—For a comprehensive understanding of disordered AMs, we examine the spectral statistics using level spacing ratio (LSR).
+Spectral statistics.—For a comprehensive understanding of disordered AMs, we examine the spectral statistics based on level spacing ratio (LSR).
 The LSR is defined as $r_{n}=\frac{\mathrm{min}(s_{n},s_{n-1})}{\mathrm{max}(s_{n},s_{n-1})}$, where $s_{n}\equiv E_{n+1}-E_{n}$ is the spacing between adjacent energy levels $E_{n}$ and $E_{n+1}$ Oganesyan and Huse (2007).
-Figure S3(a) plots disorder-averaged LSR $\langle r\rangle$ against disorder strength $W$.
+Figure S4(a) plots the disorder-averaged LSR $\langle r\rangle$ against disorder strength $W$.
 For a finite range of $W<W_{c}$, $\langle r\rangle$ remains nearly constant at $\langle r\rangle\approx 0.6$, consistent with the metallic nature of AMMMs ($\langle r\rangle_{M}=0.6$).
@@
 The probability distributions $P(r)$ at representative $W$ are shown in Fig.
-S3(b).
+S4(b).
 For weak disorder ($W=3t$), $P(r)$ closely follows the Gaussian unitary ensemble (GUE) form $P_{\text{GUE}}(r)=\frac{81\sqrt{3}}{2\pi}\frac{(r+r^{2})^{2}}{(1+r+r^{2})^{4}}$ Atas et al.
@@
 For strong disorder ($W=14t$), $P(r)$ instead matches the Poisson distribution $P_{p}(r)=\frac{2}{(1+r)^{2}}$, indicative of uncorrelated, localized states.
-In the SM, we also analyze inverse participation ratio statistics Li2 (b).
-These results further confirm the existence of a robust AMMM phase and the disorder-induced metal–insulator transition.
-Figure 3: (a) Averaged LSR $\langle r\rangle$ as a function of $W$. (b) Distribution of LSR in the metallic and insulating limits.Other parameters are as the same as Fig. 1.
+In the SM Li (2), we also analyze inverse participation ratio statistics.
+These results further confirm the existence of a robust AMMM phase and the disorder-induced metal–insulator transition  [^3] Note that there exists a mobility edge in the whole spectrum, but we focus on the properties at the Fermi energy here.
+Figure 3: (a) Averaged LSR $\langle r\rangle$ as a function of $W$.
+(b) Distribution of LSR in the metallic and insulating limits.
+We take $2\times 10^{3}$ disorder configurations.
+Other parameters are as the same as Fig.
+1.
 Phenomenological picture.—We present a phenomenological argument to interpret the KT-type phase transition in disordered AMs.
 In the 2D XY model, the KT transition arises from the unbinding of vortex-antivortex pairs as temperature increases Kosterlitz and Thouless (1973).
-Analogously, the spin-up and spin-down degrees of freedom in disordered AMs can be viewed as vortex-like and antivortex-like excitations, respectively.
-Weak magnetic disorder, described by the terms $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$, couples spin-up and spin-down states (quantized along $\sigma_{z}$), effectively forming spin-paired states ($\bm{\uparrow}\leftrightsquigarrow\bm{\downarrow}$) akin to bound vortex-antivortex pairs thermally activated in the XY model.
-Crucially, the $[C_{2}||C_{4z}]$ spin symmetry (equivalently, the $C_{4}\mathcal{T}$ symmetry that combines four-fold rotation and time-reversal symmetries) protects the marginal metallic phase Wang et al.
-(2024).
+Analogously, in 2D AMs, magnetic disorder induces spin flips that generate pronounced in-plane spin components, whose spatial texture admits vortices and antivortices Li (2).
+The disorder strength $W$ plays the role of temperature in the XY model.
+At weak $W$, we can directly visualize the appearance of in-plane spin vortex-antivortex pairs Li (2) corresponding to the marginal metallic phase.
+Note that the $C_{2}||C_{4z}]$ symmetry (equivalently, the $C_{4z}\mathcal{T}$ symmetry that combines four-fold rotation and time-reversal symmetries) can protect the marginal metallic phase Wang et al.
+(2024) [^4] The marginal metal phase can be removed by explicitly breaking $C_{4z}\mathcal{T}$, for instance, by adding a term $t_{e}(\cos k_{x}+\cos k_{y})\sigma_{z}$ to the Hamiltonian $H_{0}({\bf k})$.
 Indeed, as shown in Fig.
@@
 In absence of AM ($t_{J}=0$) or in conventional antiferromagnets, no marginal metal phase emerges, underscoring the essential role of altermagnetism.
-When the disorder strength exceeds $W_{c}$, magnetic scattering between opposite spins becomes so strong that the spin pairs are fully depaired, resembling the breakdown of vortex-antivortex pairs in the XY model.
+The marginal metallic phase corresponds to the quasi-ordered phase in the XY model at low temperatures.
+When the disorder strength exceeds $W_{c}$, magnetic scattering between opposite spins becomes so strong that the in-plane spin vortices proliferate and unpair Li (2), resembling the breakdown of vortex-antivortex pairs in the XY model at high temperatures.
 The system then enters an insulating state.
@@
 (b) The same as (a) but for $W=6t$.
+We take $5\times 10^{4}$ disorder configurations.
 Other parameters are $t_{J}=0.3t,$ $\mu=-2t$, and $\eta=10^{-3}t$.
-Mitigation of spin anisotropy by disorder.—Disorder significantly affects the characteristic spin anisotropy of AMs, namely, the alternating spin splitting in momentum space.
+Mitigation of spin anisotropy by disorder.Disorder significantly affects the characteristic spin anisotropy of AMs, namely, the alternating spin splitting in momentum space.
 To illustrate this, we track the evolution of spin anisotropy during the phase transition by analyzing spin-resolved LDOS.
@@
 $A_{\sigma}(E,{\bf k})=-\frac{1}{\pi}\mathrm{Im}[E+i\eta-\langle H({\bf k})\rangle]_{\sigma\sigma}^{-1},$ (3)
-where $\eta$ is a small spectral broadening Li2 (b).
+where $\eta$ is a small spectral broadening Li (2).
 As shown in Fig.
 4(a), the spin anisotropy remains prominent in the AMMM phase at weak disorder: spin-up and spin-down Fermi surfaces exhibit distinct orientations, confirming the persistence of spin-splitting.
-As disorder strength $W$ increases, the bands broaden due to disorder scattering, and the Fermi-surface splitting gets increasingly blurred [see Figs.
+As the disorder strength $W$ increases, the bands broaden due to disorder scattering, and the Fermi-surface splitting gets blurred [see Figs.
 4(a,b) for comparison].
@@
 (c) $G_{P}$ and $G_{AP}$ as functions of $W$ at different $E$.
-(d) TMR as a function of $W$.
-Inset: Energy spectrum of an AM ribbon of width $L_{y}=40$ along $x$ direction.
+(d) TMC as a function of $W$.
+Inset: Energy spectrum of an AM ribbon of width $L_{y}=40a$ along $x$ direction.
+We take $2\times 10^{4}$ disorder configurations.
 Other parameters are as the same as Fig.
 1.
-Tunneling magnetoresistance in disordered AMs.The spin splitting affected by disorder in AMs can also be detected via tunneling magnetoresistance in tunnel junctions.
+Tunneling magnetoconductance in disordered AMs.The spin splitting affected by disorder in AMs can also be detected via tunneling magnetoconductance in tunnel junctions.
 We consider a junction composed of two AM layers with either parallel or antiparallel Néel vectors, separated by a tunnel barrier, as illustrated in Fig.
@@
 This increase may be attributed to the fact that AM order reduces while the states are not fully localized.
-To quantify the sensitivity of tunneling to the AM bilayer configuration, we further compute the tunneling magnetoresistance $\text{TMR}=\frac{G_{P}-G_{AP}}{G_{AP}}$.
+To quantify the sensitivity of tunneling to the AM bilayer configuration, we further compute the tunneling magnetoconductance $\text{TMC}=\frac{G_{P}-G_{AP}}{G_{AP}}$.
 As shown in Fig.
-5(d), TMR depends on energy $E$.
+5(d), TMC depends on energy $E$.
 At $E=2t$, it decreases monotonically to zero, while at $E=0$, it first increases to a maximum before it eventually drops to zero.
-The latter case suggests that within a specific range of disorder strengths, the tunneling magnetoresistance can be anomalously enhanced.
+The latter case suggests that within a specific range of disorder strengths, the tunneling magnetoconductance can be anomalously enhanced.
 Conclusion and discussion.—In summary, we have shown that 2D metallic AMs persist as a marginal metal over a finite range of disorder strengths and undergo a KT metal-insulator transition as disorder increases further.
-We have also provided a phenomenological interpretation of this transition based on vortex-antivortex-like spin pairs.
-During the transition, the spin anisotropy of the AM gradually weakens, and the tunneling magnetoresistance becomes indistinguishable.
-These predictions are experimentally accessible and may help explain conflicting experiments on AM candidate materials.
+We have also provided a phenomenological interpretation of this transition based on spin vortex-antivortex pairs.
+During the transition, the spin anisotropy of the AM gradually weakens, and the tunneling magnetoconductance becomes indistinguishable.
+These predictions are experimentally accessible and may help to explain conflicting experiments on AM candidate materials.
 The key ingredient underlying our results is the interplay between AM order and magnetic disorder.
@@
 (2025), $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ with layered structures Zhang et al.
-(2025); Jiang et al.
+(2025a); Jiang et al.
 (2025), ferroelectric systems Zhu et al.
-(2025b), monolayer Ma et al.
+(2025b), monolayer materials Ma et al.
 (2021b) and twisted magnetic bilayers Liu et al.
 (2024b).
-Although our study focuses on representative $d$-wave AMs, we expect our conclusions to extend to other symmetry types, such as $g$- and $i$-wave AMs.
-We thank Lunhui Hu, K.
+Although our study focuses on representative $d$-wave AMs, the conclusions hold for other symmetry types such as $g$-wave AMs and other lattice geometries such as the Lieb lattice.
+They remain valid even when the disorder is spatially correlated Li (2).
+We thank Fakher Assaad, Lunhui Hu, K.
 Mæland, E.
-Petermann, and Jinsong Xu for helpful discussions.
+Petermann, Jinsong Xu and Yan-Yang Zhang for helpful discussions.
 This work was supported by the DFG (SFB 1170 ToCoTronics), and the Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147 (Project-Id 390858490).
+C.A.L.
+was also supported by the start-up fund at HFNL.
 H.G.
-acknowledge support from the NSFC grant No.
+acknowledges support from the NSFC Grant No.
 12074022 and the BNL-CMP open research fund under Grant No.
 2024BNL-CMPKF023.
-S.B.Z was supported by the start-up fund at HFNL, and the Innovation Program for Quantum Science and Technology (Grant No.
-2021ZD0302801).
+S.B.Z was supported by the start-up fund at HFNL, the Innovation Program for Quantum Science and Technology (Grant No.
+2021ZD0302801), and the National Natural Science Foundation of China (Grant No.
+12488101).
+Note added: Recently, we came across a related work that studies the influence of disorder on $d$-wave altermagnets with strong spin-orbit coupling and predicts a KT-type phase transition Zeng et al.
+(2025).
+However, our work substantially differs from Ref.
+Zeng et al.
+(2025) in that the KT-type transition in our case is solely induced by magnetic disorder in AMs in the absence of spin-orbit coupling and band topology.
 Supplemental materials of “Marginal Metals and Kosterlitz-Thouless Type Phase Transition in Disordered Altermagnets”
@@
 Specifically, we consider a long ribbon geometry along $x$-direction with width $L_{y}=L$ and length $L_{x}\gg L$.
-In practice, we take $L_{x}=3\times 10^{6}$.
+In practice, we take $L_{x}=3\times 10^{6}a$ in units of the lattice constant $a$.
 Periodic boundary conditions are taken along $y$-direction.
 We partition the ribbon lattice to $L_{x}$ layers along $x$-direction.
-Focusing on three consecutive layers at $x=n-1$, $n$, and $n+1$, respectively, their wave function amplitudes can be connected by
+Focusing on three consecutive layers at x=n1x=n-1, $n$, and $n+1$, respectively, their wave function amplitudes can be connected by
 $\left(\begin{array}[]{c}\psi_{n+1}\\ \psi_{n}\end{array}\right)=T_{n}\left(\begin{array}[]{c}\psi_{n+1}\\ \psi_{n}\end{array}\right),$ (S1.1)
@@
 Here, in Fig.
-S1(a), we further present the localization length for $t_{J}=2.0$ as a representative case of strong altermagnetic strength $t_{J}$.
+S1(a), we further present the localization length for $t_{J}=2t$ as a representative case of strong altermagnetic strength $t_{J}$.
 Clearly, the critical disorder strength $W_{c}$ shifts to higher values, reflecting the enhanced resilience of altermagnetic order.
@@
 S1(b)], we find $W_{c}=11.27t\pm 0.33t$.
-We find that all data converge to a single curve under this scaling formula.
-Figure S1: Kosterlitz-Thouless (KT) metal-insulator transition for $t_{J}=2.0t$.
+All data converge to a single curve under this scaling formula.
+Figure S1: Kosterlitz-Thouless (KT) metal-insulator transition for $t_{J}=2t$.
 (a) Normalized localization length $\lambda/L$ as a function of $W$ for increasing $L$.
@@
 Appendix S2 Disorder averaged effective Hamiltonian
-In this section, we describe in details how to obtain the disorder-averaged effective Hamiltonian $\langle H\rangle$.
+In this section, we describe in detail how to obtain the disorder-averaged effective Hamiltonian $\langle H\rangle$.
 Consider the altermagnetic metal on a $L\times L$ square lattice with periodic boundary conditions in both $x$ and $y$ directions.
@@
 $G_{\mathrm{avg}}^{R}({\bf r}-{\bf r}^{\prime},\omega)$ $=\langle G^{R}({\bf r},{\bf r}^{\prime};\omega)\rangle_{\mathrm{dis}}.$ (S2.2)
-It depends only on the position difference ${\bf r}-{\bf r}^{\prime}$. Next, we next Fourier transform the disorder-averaged Green’s function into momentum space,
+It depends only on the position difference ${\bf r}-{\bf r}^{\prime}$. Next, we Fourier transform the disorder-averaged Green’s function into momentum space,
 $G^{R}({\bf k},\omega)$ $=\int d({\bf r}-{\bf r}^{\prime})G_{\mathrm{avg}}^{R}({\bf r}-{\bf r}^{\prime},\omega)e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}.$ (S2.3)
-Finally, the disorder-average effective Hamiltonian is then given by
+Finally, the disorder-averaged effective Hamiltonian is then given by
 $\langle H({\bf k})\rangle=-[G^{R}({\bf k},\omega=0)]^{-1}.$ (S2.4)
-The spin-resolved density of states then follows from
+The spin-resolved density of states follows from
 $A_{\sigma}(E,{\bf k})=-\frac{1}{\pi}\mathrm{Im}[E+i\eta-\langle H({\bf k})\rangle]_{\sigma\sigma}^{-1}.$ (S2.5)
 Appendix S3 Scaling behavior under different disorder types
+In this section, we consider the scaling behavior of disordered altermagnets (AMs) under different disorder types.
 In Fig.
-S2(a), we consider on-site nonmagnetic (spin-independent) disorder, i.e.
-$w_{i}^{0}\sigma_{0}$.
-Since this type of disorder does not mix opposite spins, the altermagnets (AMs) decouple into two independent spin sectors.
+S2(a), we consider on-site nonmagnetic (spin-independent) disorder, i.e., $w_{{\bf r}}^{0}\sigma_{0}$.
+Since this type of disorder does not mix opposite spins, the AMs decouple into two independent spin sectors.
 From Fig.
 S2(a), we find that the normalized localization length always decreases with system size $L$.
-This indicates an immediate transition into insulating phases and the absence of Kosterlitz-Thouless (KT) type phase transition, consistent with conventional 2D electron behavior.
-This behavior changes completely if we consider magnetic disorder, i.e., $(w_{i}^{x}\sigma_{x}+w_{i}^{y}\sigma_{y}+w_{i}^{z}\sigma_{z})$.
-Here, the diagonal type $w_{i}^{z}\sigma_{z}$ is included for comparison, although it does not flip spins.
+This indicates an immediate transition into the insulating phase and the absence of the Kosterlitz-Thouless (KT) type phase transition, consistent with conventional 2D electron behavior.
+This behavior changes completely if we consider magnetic disorder of the type $(w_{{\bf r}}^{x}\sigma_{x}+w_{{\bf r}}^{y}\sigma_{y}+w_{{\bf r}}^{z}\sigma_{z})$.
+Here, the diagonal type $w_{{\bf r}}^{z}\sigma_{z}$ is included for comparison, although it does not flip spins.
 As shown in Fig.
-S2(b), $\Lambda=\lambda/L$ remains independent of $L$ within a wide range of disorder strengths, indicating the existence of a marginal metallic state.
-Beyond the critical value of disorder strength, $\Lambda$ decreases with $L$, indicating the KT-type metal-insulator phase transition.
+S2(b), $\Lambda=\lambda/L$ remains independent of $L$ within a wide range of disorder strengths, indicating the existence of a marginal metallic phase.
+Beyond the critical value of the disorder strength, $\Lambda$ decreases with $L$, indicating the KT-type metal-insulator transition.
 Comparing these results, we conclude that the magnetic scattering between opposite spins is essential for the occurrence of KT-type phase transitions in 2D AM metals.
@@
 Other parameters are $t_{J}=0.5t$ and $\mu=-2t$.
-Appendix S4 Spectrum statistics in disordered altermagnetism from inverse participation ratio
-In this section, we analyze the statistics from the eigenstates of the disordered AMs. To this end, we employ the inverse participation ratio (IPR) defined as (Li et al., 2017; Pixley et al., 2015)
-$I_{n}=\sum_{i,\sigma}|\psi_{n}(i,\sigma)|^{4},$ (S4.1)
+We further consider correlated disorder of Gaussian type.
+The disorder takes the form $V({\bf r})=V_{0}({\bf r})\sigma_{0}+V_{x}({\bf r})\sigma_{x}+V_{y}({\bf r})\sigma_{y}$, with zero mean $\langle V_{s}({\bf r})\rangle=0$ $(s=0,x,y)$ and spatial correlations $\langle V_{s}({\bf r})V_{s}({\bf r}^{\prime})\rangle=(W^{2}/12)\times\exp(-\left|{\bf r}-{\bf r}^{\prime}\right|^{2}/2\xi^{2})$, characterized by the correlation length $\xi$ and disorder strength $W$.
+A representative disorder configuration is shown in Fig.
+S3(a).
+To demonstrate the persistence of the marginal metallic phase and the associated KT-type phase transition, we calculate the scaling of the conductance as a function of $W$, see Fig.
+S3(b).
+The conductance curves merge together for different system sizes $L$ in the marginal metallic phase (for $W<W_{c}$) while they decrease with increasing $L$ in the insulating phase (for $W>W_{c}$) after a critical disorder strength $W_{c}$.
+Figure S3: (a) A disorder configuration of the Gaussian-correlated disorder in real space with $\xi=1$ and $W=2t$.
+(b) Averaged conductance as a function of disorder strength $W$ for the correlated disorder.
+We take $2\times 10^{4}$ disorder configurations in the conductance scaling.
+Other parameters are $t_{J}=0.3t$ and $\mu=-2t$.
+Appendix S4 Spectrum statistics from inverse participation ratio
+In this section, we analyze the statistics of the eigenstates of the disordered AMs. To this end, we employ the inverse participation ratio (IPR) defined as (Li et al., 2017; Pixley et al., 2015)
+$I_{n}=\sum_{{\bf r},\sigma}|\psi_{n}({\bf r},\sigma)|^{4},$ (S4.1)
 where the summation runs over all sites and spins.
 The disorder-averaged IPR $\langle I\rangle$ takes nearly zero values at weak disorder strength, consistent with the altermagnetic marginal metals discussed in the main text [Fig.
-S3(a)].
-It becomes finite when the system enters an insulating phase.
+S4(a)].
+It becomes finite when the system enters the insulating phase.
 The IPR scaling defines a fractal dimension $d_{2}$ via $\langle I\rangle\propto L^{-d_{2}}$.
 From Fig.
-S3(b), we find $d_{2}=1.84$ in the altermagnetic marginal metals, which is close to the value $d_{2}^{M}=2$ for ideal metals.
+S4(b), we find $d_{2}=1.84$ in the marginal metal, which is close to the value $d_{2}^{M}=2$ for ideal metals.
 For the insulating state at $W=14t$, we find $d_{2}=0.14$, consistent with the value $d_{2}^{I}=0$ for insulators.
@@
 As shown in Fig.
-S3(c), we find $\chi=0.035$ at $W=3t$ and $\chi=0.44$ at $W=14t$, both in excellent agreement with the values inferred from $d_{2}$.
-Thus, the conjectured relation is confirmed in both the altermagnetic marginal metals and insulating states.
+S4(c), we find $\chi=0.035$ at $W=3t$ and $\chi=0.44$ at $W=14t$, both in excellent agreement with the values inferred from $d_{2}$.
+Thus, the conjectured relation is confirmed in both the marginal metallic and insulating phases.
 Figure S4: (a) Disorder-averaged IPR as a function of disorder strength $W$.
@@
 Other parameters for all plots are $t_{J}=0.3t$, and $\mu=-2t$.
+Appendix S5 $d$-wave altermagnets on the Lieb lattice
+In this section, we consider $d$-wave AMs on a different lattice geometry, i.e., the Lieb lattice.
+There are three distinct sites in each unit cell, labeled as $A,B$, and $C$, respectively, as shown in Fig.
+S5(a).
+The $d$-wave AMs can be described by the Hamiltonian (Brekke et al., 2023)
+$H_{\mathrm{Lieb}}=t\sum_{\langle ij\rangle\sigma}C_{i,\sigma}^{\dagger}C_{j,\sigma}-J\sum_{i,\sigma,\sigma^{\prime}}C_{i,\sigma}^{\dagger}{\bf S}_{i}\cdot{\bf\bm{\sigma}}_{\sigma\sigma^{\prime}}C_{i,\sigma^{\prime}}-\mu\sum_{i,\sigma}C_{i,\sigma}^{\dagger}C_{i,\sigma}+\varepsilon_{B}\sum_{i\in B,\sigma}C_{i,\sigma}^{\dagger}C_{i,\sigma},$ (S5.1)
+where $t$ is the hopping amplitude, $J$ describes the exchange interaction strength between itinerant electrons and local spins ${\bf S}_{i}$, $\mu$ is the chemical potential, and $\varepsilon_{B}$ is the on-site energy at the nonmagnetic sites.
+Note that there is no spin-orbit coupling, thus the Hamiltonian is block diagonal in spin space $\sigma$.
+The Bloch Hamiltonian in momentum space for spin $\sigma$ reads
+$H_{\sigma}({\bf k})=\left(\begin{array}[]{ccc}-\sigma JS-\mu&t(1+e^{ik_{x}})&0\\ t(1+e^{-ik_{x}})&-\mu+\varepsilon_{B}&t(1+e^{ik_{y}})\\ 0&t(1+e^{-ik_{y}})&+\sigma JS-\mu\end{array}\right).$ (S5.2)
+We plot the energy spectrum along high-symmetry lines in Fig. S5(b). The energy bands exhibit anisotropic spin splitting, a characteristic feature of $d$-wave AMs.
+Our results of disordered AMs can be generalized to the Lieb lattice.
+To illustrate this, we consider the same disorder of the form $(w_{i}^{0}\sigma_{0}+w_{i}^{x}\sigma_{x}+w_{i}^{y}\sigma_{y})$ at each lattice site and perform the scaling of the conductance of the $d$-wave AM on the Lieb lattice.
+We find that the conductance curves for different system sizes $L$ collapse when the disorder strength $W$ is below a critical value $W_{c}\simeq 3t$, signaling the emergence of a marginal metallic phase.
+For large $W$ above $W_{c}$, the conductance decreases with increasing $L$, consistent with an insulating phase.
+These results demonstrate the persistence of the marginal metallic phase and KT-type phase transition in the disordered $d$-wave AMs on the Lieb lattice.
+Figure S5: (a) Altermagnet in a two-dimensional Lieb lattice, which contains three distinct sites in each unit cell, labeled by blue, gray, and red color respectively.
+(b) The electron bands of AMs on Lieb lattice along the high symmetry lines.
+(c) Averaged conductance as a function of disorder strength $W$.
+We average over $2\times 10^{3}$ disorder configurations.
+Other parameters for panels (b) and (c) are $JS=0.6t$, $\varepsilon_{B}=0$, and $\mu=-2t$.
+Appendix S6 $g$-wave altermagnets and KT-type phase transition
+In this section, we show that our results on $d$-wave AMs can be generalized to $g$-wave AMs in 2D. For illustration, we consider the minimum effective Hamiltonian for $g$-wave AMs (Ezawa, 2025)
+$H_{g}({\bf k})=(-\mu+2t-t\cos k_{x}-t\cos k_{y})\sigma_{0}+J\sin k_{x}\sin k_{y}(\cos k_{x}-\cos k_{y})\sigma_{z},$ (S6.1)
+where $t$ is the usual hopping amplitude, $J$ is the altermagnetic strength, $\mu$ is chemical potential, and $\sigma_{0}$ and $\sigma_{z}$ are the identity and Pauli matrices in spin space, respectively.
+The Fermi surface is shown in Fig.
+S6(a).
+It clearly exhibits the g-wave form of spin splitting.
+We consider the same onsite disorder of the form $(w_{{\bf r}}^{0}\sigma_{0}+w_{{\bf r}}^{y}\sigma_{y}+w_{{\bf r}}^{z}\sigma_{z})$ as in the main text.
+We perform the scaling calculation of the conductance for the $g$-wave AM, as shown in Fig.
+S6(b).
+The conductance curves converge for different system sizes $L$ when the disorder is smaller than a critical values $W_{c}\simeq 3.4t$, indicating the appearance of a marginal metallic phase.
+While the conductance decreases as the system size $L$ grows for disorder strength larger than $W_{c}$.
+These results demonstrate the presence of the marginal metallic phase and KT-type phase transition in the disordered $g$-wave AMs.
+Figure S6: (a) Fermi surface of the $g$-wave AM at the Fermi energy, where the spin-up and spin-down states are indicated by different color and arrows.
+(b) Averaged conductance as a function of disorder strength $W$.
+We average over $2\times 10^{3}$ disorder configurations in the conductance scaling.
+Other parameters are $J=1.2t$ and $\mu=-0.6t$.
+Appendix S7 Phenomenological picture based on spin vortices
+In this section, we provide more details to support the phenomenological picture based on spin vortices in disordered AMs.
+We first briefly review the mechanism responsible for the KT phase transition.
+In the 2D XY model (and thin-film superfluids), the KT phase transition happens due to the formation of vortex-antivortex pairs (Kosterlitz and Thouless, 1973).
+At zero temperature, there are no vortices in the system.
+At low temperatures, bounded vortex-antivortex pairs are thermally activated.
+In this low-temperature regime, the system is in a quasi-ordered phase for the XY model and a superfluid phase for thin-film superfluids.
+As the temperature increases, the numbder of vortices increases and vortex-antivortex pairs become unbounded after a critical temperature, leading to the KT phase transition.
+At high temperatures above the transition, the system is in a disordered phase for the XY model and a normal fluid phase for thin-film superfluid.
+The disorder-induced KT-type phase transition in 2D AMs can be explained along similar lines, with the disorder strength $W$ playing the role of temperature in the XY model or thin-film superfluid.
+In the 2D AMs with strongly anisotropic spin splitting, magnetic disorder of the form $(w_{{\bf r}}^{0}\sigma_{0}+w_{{\bf r}}^{x}\sigma_{x}+w_{{\bf r}}^{y}\sigma_{y})$ induces spin flips and generate pronounced in-plane magnetization components whose spatial texture forms a spin texture admitting vortices and antivortices (phase windings $\text{$\theta({\bf r})$=$\pm$}2\pi$).
+The typical in-plane spin textures for a vortex and anitvortex are schematically shown in Fig.
+S7(a).
+To substantiate this, we calculate numerically the real-space-resolved in-plane spin components at the Fermi level in presence of magnetic disorder:
+$\bar{\sigma}_{s}({\bf r})=\sum_{n}\langle\Psi_{n}({\bf r})\left|\sigma_{s}\right|\Psi_{n}({\bf r})\rangle\rho(E_{n}),$ (S7.1)
+where $\Psi_{n}({\bf r})$ is the eigenstate, $E_{n}$ the eigenvalue, and $\rho(E_{n})$ the density of states $\rho(E_{n})=\frac{\Gamma}{E_{n}^{2}+\Gamma^{2}}$ with $\Gamma$ a small band broadening.
+We illustrate the results in Fig.
+S7(b-d).
+At $W=0$, there are no vortices in the system, just as in the 2D XY model and thin-film superfluid at $T=0$.
+Remarkably, in the weak-disorder regime, we observe bound vortex-antivortex pairs appear in the in-plane spin texture [see Fig.
+S7(b)].
+As the disorder strength increases, the number of such pairs grows [see Figs.
+S7(c) and S7(d)], similar to thermally activated vortex pairs in the 2D XY model and thin-film superfluid.
+In this case, the system is in a marginal metallic phase, corresponding to the quasi-ordered phase in XY model or the superfluid phase in thin-film superfluids at low temperatures.
+In the strong-disorder regime, vortices significantly proliferate and unbind, that is to say, the phase field $\theta({\bf r})$ gets more and more distorted and in-plane spin correlations become short-ranged [see Fig.
+S7(d)].
+In this case, the system is an insulator, corresponding to the disordered phase in the XY model or the normal fluid phase in thin-film superfluids at high temperatures.
+This progression, from bound pairs to vortex proliferation and vortex unbinding, provides a consistent, microscopic picture that underscores the KT transition induced by magnetic disorder in AMs.
+Note that no vortices appear for purely nonmagnetic disorder of $w_{{\bf r}}^{0}\sigma_{0}$ type or spin-polarized disorder of only $w_{{\bf r}}^{x}\sigma_{x}$ type and only $w_{{\bf r}}^{y}\sigma_{y}$ type.
+Accordingly, no KT-type phase transition occurs under these types of disorder.
+This further strengthens our physical picture based on the formation of in-plane spin vortices.
+Figure S7: Formation of vortex and antivortex in the in-plane magnetization.
+(a) Sketch of vortex and antivortex in the in-plane spin components.
+(b) Spin texture at delocalized/marginal metal regime with disorder strength $W=0.5t$.
+Blue and yellow spots mark vortex and antivortex cores.
+(c) Same as (b) but at localized regime with $W=5t$, showing the proliferation of unbounded vortices and antivortices.
+(d) Averaged number of vortex pairs as a function of disorder strength $W$.
+Here, 50 disorder configurations are taken into the average.
+We use the parameters: $t_{J}=0.3t$, $\mu=-2t$, $\xi=2$, and $\Gamma=0.05t$ for illustration.